OpenStudy (anonymous):

A company will need $70,000 in 5 years for a new addition. To meet this goal, the company deposits money in a account today that pays 7% annual interest compounded quaterly. Find the amount that should be invested to total $70,000 in 5 years?

OpenStudy (kropot72):

The equation to solve this is as follows: \[70,000=P(1+(0.07\div3))^{5\times 3}=P(1.0233333)^{15}= P \times 1.41337\] Therefore P = 70,000/1.41337 = $49,527 Where P is the amount that should be invested.